OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form

OpenStudy (anonymous):

Y = mx + b ( passing through -3,4 ) 4 = 2( -3) +b 4 = -6 + b(adding 6 to both sides) 6 = 6 10= b y = 2x + 10 - slope intercept form y - ( 4) = 2 ( x- -3) y-4= 2x +6 ( point slope form)